OpenStudy (anonymous):

13.6 Determine the following derivatives with help of the chain rule, maybe applied several times. In each case, write down the inner and outer functions used. a) \(\Large h(x) = cos(sin x) \)

OpenStudy (experimentx):

?? what inner and outer functions

OpenStudy (anonymous):

i am not sure what it is but i asked before here this kind of question there were people who was familiar with it

OpenStudy (experimentx):

f(x) = cos(x) g(x) = sin(x) h(x) = f(g(x)) perhaps

OpenStudy (anonymous):

ok to start with it good for me maybe its right i ask some more people to be sure.. thanks

OpenStudy (experimentx):

np :)

OpenStudy (anonymous):

\[-\sin (\sin(x))*\cos(x)\] Where -sin(sin(x) is the outer function in respect to the inner function and cos(x) is the inner function

OpenStudy (anonymous):

@mahmit2012 what you think about Spacelimbus's solution?

OpenStudy (anonymous):

thank you Spacelimbus :) i dont know too what the solution is, lets wait.. what it comes out

OpenStudy (anonymous):

h(x) =cos(sinx) , i think the inner f is sinx and the outer f is cos. now by chain rule, let u=sinx , then du/dx = cosx therefore cos(sinx) becomes cosu, and dy/du =-sinu then dy/dx =dy/du *du/dx,just substute the values. nb h(x) = y

OpenStudy (anonymous):

ok thank you pota by the way your son - daughter nice sweet baby

OpenStudy (anonymous):

Here you have the solution again computed by WolframAlpha. With Leibniz notation it is easier to see what you are doing in this case. \[h(x)=\cos(\sin(x))\] now let u=sin(x) then \[du/dx = \cos(x)\] and \[dy/du= -\sin(x)\] so that \[du/dx*dy/du= dy/dx = -\sin(u)\cos(x) = -\sin(\sin(x)\cos(x)\]

OpenStudy (anonymous):

thank you very much Spacelimbus it looks good

OpenStudy (anonymous):

you are welcome